# Transformer Block Coupling and its Correlation with Generalization in LLMs

### v3

@@ 6.1 Emergence of Regularity with Coupling
Under certain assumptions, the emergence of increased linearity and exponential spacing in many LLMs can be analyzed as a result of coupling. Considering input embeddings $x_{1}^{0},\ldots,x_{n}^{0}$ and the linearization of the last token embedding $x_{n}^{l}$ given by $J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})$:
$x_{n}^{l}=(I+J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0}))x_{n}^{l-1}$
We simplify notation and write $x_{n}^{l}=x^{l},\,J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})=J^{l}$. The representations follow the linearized equation:
@@
Assuming that $J_{l}\approx USU^{T}$, this equation predicts that the norm of $x_{l}$ would exhibit exponential growth layer by layer. Expanding $U$ and $S$,
@@
where $u_{j}$ and $s_{j}$ represent the eigenvectors and eigenvalues of the Jacobian, respectively. In general, trajectories are not expected to be perfectly linear unless $x^{0}$ aligns with an eigenvector of $J$. However, in our experiments, we observe a notable tendency towards linearity, suggesting that the representations align progressively during training with the eigenvectors of the coupled Jacobians.
6.2 Significance of Coupling
The emergence of coupling with training steps (Figure 1) may provide insight into the dynamics. Under full coupling and a difference equation approximation, the representations evolve as
$x^{l}=\sum_{j=1}^{d_{\text{model}}}u_{j}(1+s_{j})^{l}u_{j}^{\top}x^{0}$
Due to the coupling, the dynamics exhibit either exponential growth or decay in different subspaces, depending on the sign of $s_{j}$, which is known to cause challenges for optimization as in past works on dynamical isometry (Pennington et al., 2017). We infer that increasing coupling during training makes optimization progressively more difficult. Conversely, as training progresses, it becomes harder to achieve stronger coupling, and is consistent with the logarithmic trend in Figure 1.
@@ 7 Related Work
Residual Networks. ResNets (He et al., 2016) have been viewed as an ensemble of shallow networks (Veit et al., 2016), with studies delving into the scaling behaviour of their trained weights (Cohen et al., 2021). The linearization of residual blocks by their Residual Jacobians was first explored by Rothauge et al. (2019), who examined Residual Jacobians and their spectra in the context of stability analysis, and later by Li & Papyan (2023) who discovered Residual Alignment. Coupling of Jacobian singular vectors in LLM transformers extends previous results for classifier Resnets (Li & Papyan, 2023). We show coupling in $J_{l}^{t_{1}t_{2}}$ across various tokens, which is specific to tokenization in transformers, in addition to demonstrating coupling of $J_{l}^{t_{1}t_{2}}$ across $l$, which was also identified analogously in ResNets (Li & Papyan, 2023). Further comparison is included in Appendix A.3.
Neural Ordinary Differential Equations. Neural ODEs (Chen et al., 2018) view ResNets as a discretized dynamical process, with past work (Sander et al., 2022) showing the convergence of Residual Networks to a system of linear ODE, with some extensions to transformers (Zhong et al., 2022; Li et al., 2021) The emergence of coupling in transformers suggests that a discretization of a simple itertaive process emerges in LLMs.
@@
Hidden Representation Dynamics. Prior research interprets deep neural networks through dynamical systems, revealing that training trajectories align with geodesic curves (Gai & Zhang, 2021) and partition activation space into basins of attraction (Nam et al., 2023). For further related works, see (Geshkovski et al., 2023b; a; Tarzanagh et al., 2023; Valeriani et al., 2023).
Structure in Hidden Representations. Neural Collapse (Papyan et al., 2020) highlights emergent regularity in last-layer representations, with subsequent studies exploring hidden-layer structures and their theoretical underpinnings (Wang et al., 2024a; Parker et al., 2023; Zangrando et al., 2024; Garrod & Keating, 2024; Wang et al., 2024b; Hoyt & Owen, 2021; Arous et al., 2023; Zarka et al., 2021; Ben-Shaul & Dekel, 2022; Papyan, 2020; Súkeník et al., 2023). In LLMs, recent works identify uniform token structures (Wu & Papyan, 2024) and low-dimensional hidden trajectories (Sarfati et al., 2024). Our work examines local token interactions through Jacobian dynamics across all LLM layers.
8 Limitations
Our analysis is limited to pretrained LLMs and their fine-tuned variants due to the high computational cost of training large models. Variations in experimental setups across independently trained models hinder direct comparisons, making it challenging to pinpoint causes of differing regularity. However, the consistent emergence of these properties warrants further study.
@@ 9 Conclusion
Our primary goal was to contribute to the understanding of the mechanics underlying transformer architectures through an analysis of the trajectories of token embeddings and their interactions. Our research builds on the understanding of transformer architectures by revealing the coupling, across depth and token, of singular vectors in the Jacobians of transformer blocks for multiple LLMs trained by various organizations. We establish a correlation between the strength of this coupling and benchmark performance on the HuggingFace Open LLM Leaderboard, highlighting the significance of transformer block coupling for generalization. These findings open avenues for future research, encouraging deeper exploration into the connections between regularity of hidden representations, model specifications, and generalization.
@@ A.1 Suite of Large Language Models and Prompt Data
We evaluate 38 total LLMs (24 base LLMs and 14 fine-tuned) on 6 datasets from the HuggingFace Open LLM leaderboard (Beeching et al., 2023).
@@
Falcon (40B, 7B) (Almazrouei et al., 2023), Llama-3 (70B, 8B) (AI@Meta, 2024), Llama-2 (70B, 13B, 7B) (Touvron et al., 2023), MPT (30B, 7B) (Team, 2023a; b), Mistral v0.1 (Jiang et al., 2023), Gemma (7B, 2B) (Team et al., 2024), Gemma 1.1 (7B, 2B), NeoX (20B) (Black et al., 2022), Neo (2.7B) (Black et al., 2021; Gao et al., 2020), Pythia (6.9B).(Biderman et al., 2023), and GPT-2 (1.5B, 774M, 355M, 117M) (Radford et al., 2019).
@@
CodeLlama (34B, 13B, 7B) (Rozière et al., 2024), CodeLlama Instruct (34B, 13B, 7B) (Rozière et al., 2024), Mistral-v0.1 Instruct (7.3B) (Jiang et al., 2023), Mistral-v0.2 Instruct (Jiang et al., 2023), CodeGemma (Team et al., 2024).
@@
7 B 32 4096 1 T 56.83 ALiBi
Phi 2 B 32 2560 1.4 T 61.33 RoPE
@@
1 B 24 2048 54 B RoPE
Mistral-v0.1 7.3 B 32 4096 60.97 ALiBi
Mistral-v0.1 (IT) 7.3 B 32 4096 54.96 ALiBi
Mistral-v0.2 (IT) 7.3 B 32 4096 65.71 ALiBi
@@ 1 B 24 2048 54 B RoPE
Gemma-1.1 7 B 28 3072 6 T 30.0 RoPE
2 B 18 2048 6 T 60.09 RoPE
@@
2.7 B 32 2560 0.42 36.20 Sine
Pythia 12 B 36 5120 0.3 T 58.9 RoPE
@@
2.8 B 32 2560 0.3 T RoPE
1.4 B 24 2048 0.3 T RoPE
1 B 16 2048 0.3 T RoPE
@@ A.5 Dynamical Motivation
where all the $A$ matrices are aligned. One benefit of this interpretation is that, we can write $x^{l}$ in a simple closed form, as above. We quantify the similarity of hidden trajectories to the evolution of the above difference equation, in order to detect the emergence of a simple linearization to representations. The emergence of increased linearity and exponential spacing in many LLMs can be analyzed as a result of coupling under some conditions on the spectral decomposition of the Jacobians. Considering input embeddings $x_{1}^{0},\ldots,x_{n}^{0}$ and the linearization of the last token embedding $x_{n}^{l}$ given by $J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})$:
$x_{n}^{l}=(I+J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0}))x_{n}^{l-1}$ (23)
@@
Suppose that $x^{0}=u_{k}$ is the $k$-th left singular vector of $J^{l}$. It follows that $x^{L}=\prod_{l=1}^{L}(1+s_{k}^{l})x^{0}$, where $s_{k}^{l}$ denotes the $k$-th singular value at layer $l$. The exponential spacing measurement is motivated by the consistent choice $s_{k}=s_{k}^{1}=s_{k}^{2}=\cdots=s_{k}^{L}$. Explicitly, $x^{L}=(1+s_{k})^{L}x^{0}$, and by Equation 17, for each $l$
$\alpha_{k}^{l}=\ln\left(\frac{(1+s_{k})^{l}||u_{0}||}{(1+s_{k})^{l-1}||u_{0}||}\right)=\ln(1+s_{k})\implies\text{ED}=0$
@@ A.6 LLM Evaluation and Implementation Details
Due to the high memory requirement for computing Jacobians, for experiments involving the Jacobians, NVIDIA Quadro RTX 6000 was used additionally for $13B>n_{\text{param}}\geq 7B$ and corresponding models were quantized. Additionally, due to compute restrictions, alignment of singular vectors of Jacobians was computed on a smaller subset of the 1,200 prompts.
@@ A.7 ViT Training Details
For further investigation of coupling in transformers, we train 64 Vision Transformers (ViTs) following the default configurations of DEiT training (Touvron et al., 2021) on CIFAR10 (Krizhevsky, 2009). For a fixed ViT architecture with embedding dimension 192, depth of 12 layers, and 3 attention heads, we vary the weight decay $\{0.005,0.01,0.02,0.04,0.06,0.08,0.1,0.2\}$ and stochastic depth rate $\{0,0.025,0.05,0.075,0.1,0.15,0.2,0.3\}$. Optimization uses ADAM optimizer with $45e-$ 4 learning rate and cosine scheduler for 500 epochs with 30 epochs of linear warmup. Training proceeds with data mixup using $\alpha=0.8$ Optimization proceeds on 4 NVIDIA Tesla T4 GPUs with 128 batch size and data parallelization, with total training time being approximately 2 hours.
@@ A.8 Additional Experimental Results
Figure 8: Exponential fit between coupling and performance.
Figure 9: Plotting (1) the number of tokens used to train each model against its mean depth-wise coupling score against and (2) the number of training tokens against its LLM Huggingface Benchmark score.
Figure 10: Utilizing the Easy2Hard Benchmark dataset Ding et al. (2024), the difficulty score of a prompt is plotted against the mean depth-wise coupling over that prompt. Additionally, plotted on the right is the prompt length (in tokens) plotted against the mean depth-wise coupling on that prompt.
Table 3: $R^{2}$ for Various Models (across different datasets)
Model $\bm{R^{2}}$
Llama2-7B 0.5484
MPT-30B 0.9629
Gemma-7B 0.7832
Mistral-v0.1 7B 0.5876
Subset $R^{2}$ $p$-value
All models $0.75$ $1.56\times 10^{-6}$
$7$B models only $0.55$ $0.023$
Score $>45$ $0.39$ $0.023$
Table 4: Summary of $R^{2}$ and $p$-values for different subsets of models for the correlation between coupling and performance.
@@
Figure 12: Transformer Block Coupling for Fixed Weight Decay. Coupling plotted against accuracy for 64 ViTs with stochastic depths $\{0,0.025,0.05,0.075,0.1,0.15,0.2,0.3\}$.

### v4

@@ 6.1 Emergence of Regularity with Coupling
Under certain assumptions, the increased linearity and exponential spacing observed in many models can be analyzed as a consequence of coupling. Considering input embeddings $x_{1}^{0},\ldots,x_{n}^{0}$ and the linearization of the last token embedding $x_{n}^{l}$ given by $J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})$:
$x_{n}^{l}=(I+J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0}))x_{n}^{l-1}.$
To simplify notation, we write $x_{n}^{l}=x^{l},\,J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})=J^{l}$. The representations follow the linearized equation:
@@
Assuming perfect coupling of singular vectors across depth (and between left and right singular vectors), we have $J_{l}\approx US_{l}U^{T}$. As discussed in Appendix A.5, assuming the singular values across depth are approximately constant ($S_{l}=S$), this equation predicts that the norm of $x_{l}$ would exhibit exponential growth if $x_{0}$ is some eigenvector of $J$. Expanding $U$ and $S$,
@@
where $u_{j}$ and $s_{j}$ represent the singular vectors and singular values of the Jacobian, respectively. In general, trajectories are not expected to be perfectly linear unless $x^{0}$ aligns with an eigenvector of $J$. However, in our experiments, we observe a notable tendency towards linearity, suggesting that the representations align progressively during training with the eigenvectors of the coupled Jacobians.
@@ 7 Related Work
Residual Networks. ResNets (He et al., 2016) have been viewed as an ensemble of shallow networks (Veit et al., 2016), with studies delving into the scaling behaviour of their trained weights (Cohen et al., 2021). The linearization of residual blocks by their Residual Jacobians was first explored by Rothauge et al. (2019), who examined Residual Jacobians and their spectra in the context of stability analysis, and later by Li & Papyan (2023) who discovered Residual Alignment. Coupling of Jacobian singular vectors in LLM transformers extends previous results for Resnets (Li & Papyan, 2023). We show coupling in $J_{l}^{t_{1}t_{2}}$ across various tokens, which is specific to tokenization in transformers, in addition to demonstrating coupling of $J_{l}^{t_{1}t_{2}}$ across $l$, which was also identified analogously in ResNets (Li & Papyan, 2023). Further comparison is included in Appendix A.3.
Neural Ordinary Differential Equations. Neural ODEs (Chen et al., 2018) view ResNets as a discretized dynamical process, with past work (Sander et al., 2022) showing the convergence of Residual Networks to a system of linear ODE, with some extensions to transformers (Zhong et al., 2022; Li et al., 2021) The emergence of coupling in transformers suggests that a discretization of a simple iterative process emerges in LLMs.
@@
Hidden Representation Dynamics. Prior research interprets deep neural networks through dynamical systems, revealing that training trajectories align with geodesic curves (Gai & Zhang, 2021) and partition activation space into basins of attraction (Nam et al., 2023). For further related works, see (Geshkovski et al., 2023b; a; Tarzanagh et al., 2023; Valeriani et al., 2023; Hosseini & Fedorenko, 2023).
Structure in Hidden Representations. Neural Collapse (Papyan et al., 2020) highlights emergent regularity in last-layer representations, with subsequent studies exploring hidden-layer structures and their theoretical underpinnings (Wang et al., 2024a; Parker et al., 2023; Zangrando et al., 2024; Garrod & Keating, 2024; Wang et al., 2024b; Hoyt & Owen, 2021; Arous et al., 2023; Zarka et al., 2021; Ben-Shaul & Dekel, 2022; Papyan, 2020; Súkeník et al., 2023). In LLMs, recent works identify uniform token structures (Wu & Papyan, 2024) and low-dimensional hidden trajectories (Sarfati et al., 2024). Our work examines local token interactions through Jacobian dynamics across all layers.
@@ 8 Conclusion
Our primary goal was to contribute to the understanding of the mechanics underlying transformer architectures through an analysis of the trajectories of token embeddings and their interactions. Our research builds on the understanding of transformer architectures by revealing the coupling, across depth and token, of singular vectors in the Jacobians of transformer blocks for multiple LLMs trained by various organizations, as well as ViTs. We establish a correlation between the strength of this coupling and benchmark performance on the HuggingFace Open LLM Leaderboard, highlighting the significance of transformer block coupling for generalization. These findings open avenues for future research, encouraging deeper exploration into the connections between regularity of hidden representations, training methods, and generalization.
Acknowledgments
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). This research was supported in part by the Province of Ontario, the Government of Canada through CIFAR, and industry sponsors of the Vector Institute (www.vectorinstitute.ai/partnerships/current-partners/). This research was also supported in part by Compute Ontario (https://www.computeontario.ca/) and the Digital Research Alliance of Canada (https://alliancecan.ca/).
@@ A.1 Suite of Large Language Models and Prompt Data
We evaluate a comprehensive collection of LLMs on 6 datasets from the HuggingFace Open LLM leaderboard (Beeching et al., 2023).
@@
Falcon (40B, 7B) (Almazrouei et al., 2023), Falcon2 (11B) (Malartic et al., 2024), Llama-3 (70B, 8B) (AI@Meta, 2024), Llama-2 (70B, 13B, 7B) (Touvron et al., 2023), MPT (30B, 7B) (Team, 2023a; b), Mistral v0.1 (7B) (Jiang et al., 2023), Gemma (7B, 2B) (Team et al., 2024), Gemma 1.1 (7B, 2B), Phi (1B, 1.5B, 2B) (Gunasekar et al., 2023; Li et al., 2023), OLMo (1B, 7B) (Groeneveld et al., 2024), Qwen (Bai et al., 2023; Yang et al., 2024), NeoX (20B) (Black et al., 2022), Neo (2.7B) (Black et al., 2021; Gao et al., 2020), Pythia (1.4B, 2.8B, 6.9B, 12B) (Biderman et al., 2023), and GPT-2 (1.5B, 774M, 355M, 117M) (Radford et al., 2019).
@@
CodeLlama (34B, 13B, 7B) (Rozière et al., 2024), CodeLlama Instruct (34B, 13B, 7B) (Rozière et al., 2024), Mistral-v0.1 Instruct (7B) (Jiang et al., 2023), Mistral-v0.2 Instruct (Jiang et al., 2023), CodeGemma (Team et al., 2024).
@@
Falcon2 11 B 60 4096 64.28 RoPE
Mistral-v0.1 7 B 32 4096 60.97 ALiBi
Mistral-v0.1 (IT) 7 B 32 4096 54.96 ALiBi
Mistral-v0.2 (IT) 7 B 32 4096 65.71 ALiBi
Mistral-v0.3 7 B 32 4096 60.28 ALiBi
Vicuna 7 B 32 4096 52.06 RoPE
@@
7 B 32 4096 1 T 56.84 ALiBi
OLMO 7 B 32 4096 2.5 T 43.36 RoPE
1 B 16 2048 3 T 36.78 RoPE
OLMO-1.7 7 B 32 4096 2.5 T 52.82 RoPE
Phi-2 2 B 32 2560 1.4 T 61.33 RoPE
@@
1 B 24 2048 54 B 47.69 RoPE
@@
Gemma-1.1 7 B 28 3072 6 T 60.09 RoPE
2 B 18 2048 6 T 30.0 RoPE
@@
Qwen1.5 14 B 40 5120 66.7 RoPE
7 B 32 4096 61.76 RoPE
4 B 40 2560 57.05 RoPE
1.8 B 24 2048 46.55 RoPE
0.5 B 24 1024 38.62 RoPE
Qwen2 7 B 28 3584 68.4 RoPE
@@
2.7 B 32 2560 0.42 T 36.20 Sine
Pythia 12 B 36 5120 0.3 T 38.82 RoPE
@@
2.8 B 32 2560 0.3 T 37.09 RoPE
1.4 B 24 2048 0.3 T 34.75 RoPE
1 B 16 2048 0.3 T 32.78 RoPE
@@ A.4 Significance of Coupling
The emergence of coupling with training steps (Figure LABEL:fig:coupling-mega-plot) may provide insight into the dynamics. Under full coupling and a difference equation approximation, the representations evolve as
$x^{l}=\sum_{j=1}^{d_{\text{model}}}u_{j}(1+s_{j})^{l}u_{j}^{\top}x^{0}$
Due to the coupling, the dynamics exhibit either exponential growth or decay in different subspaces, depending on whether $s_{j}$ is greater or less than $1$, which is known to cause challenges for optimization as in past work on dynamical isometry (Pennington et al., 2017). We infer that increasing coupling during training makes optimization progressively more difficult. Conversely, as training progresses, it becomes harder to achieve stronger coupling, and is consistent with the logarithmic trend in Figure LABEL:fig:coupling-mega-plot.
@@ A.5 Dynamical Motivation
where all the $A$ matrices are aligned. One benefit of this interpretation is that, we can write $x^{l}$ in a simple closed form, as above. We quantify the similarity of hidden trajectories to the evolution of the above difference equation, in order to detect the emergence of a simple linearization to representations. The emergence of increased linearity and exponential spacing in many LLMs can be analyzed as a result of coupling under some conditions on the spectral decomposition of the Jacobians. Consider input embeddings $x_{1}^{0},\ldots,x_{n}^{0}$ and the linearization of the last token embedding $x_{n}^{l}$ given by $J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})$:
$x_{n}^{l}=(I+J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0}))x_{n}^{l-1}.$ (18)
@@
Suppose that $x^{0}=u_{k}$ is the $k$-th left singular vector of $J^{l}$. It follows that $x^{L}=\prod_{l=1}^{L}(1+s_{k}^{l})x^{0}$, where $s_{k}^{l}$ denotes the $k$-th singular value at layer $l$. The exponential spacing measurement is motivated by the consistent choice $s_{k}=s_{k}^{1}=s_{k}^{2}=\cdots=s_{k}^{L}$. Explicitly, under such an assumption, $x^{L}=(1+s_{k})^{L}x^{0}$, and by Equation 12, for each $l$
$\alpha_{k}^{l}=\ln\left(\frac{(1+s_{k})^{l}||u_{0}||}{(1+s_{k})^{l-1}||u_{0}||}\right)=\ln(1+s_{k})\implies\text{ED}=0,$
@@ A.6 LLM Evaluation and Implementation Details
Due to the high memory requirement for computing block Jacobians, for experiments involving the Jacobians, NVIDIA Quadro RTX 6000 was used additionally for $13B>n_{\text{param}}\geq 7B$ and corresponding models were quantized. Additionally, coupling of singular vectors of Jacobians was computed on a smaller subset of the dataset prompts.
@@ A.7 ViT Training Details
For further investigation of coupling in transformers, we train 3 Vision Transformers (ViTs) following the default configurations of DEiT training (Touvron et al., 2021) on CIFAR10 (Krizhevsky, 2009). For a fixed ViT architecture with embedding dimension 192, depth of 24 layers, and 3 attention heads, we vary the weight decay $0.05$ and stochastic depth rate $\{0.025,0.05,0.075\}$. Optimization uses ADAM optimizer with $35e-$ 3 learning rate and cosine scheduler for 600 epochs with 10 epochs of linear warmup. Training proceeds with data mixup using $\alpha=0.8$ Optimization proceeds on 4 NVIDIA Tesla T4 GPUs with 128 batch size and data parallelization, with total training time being approximately 6 hours.
@@ A.8 Additional Experimental Results
Figure 33: Plotting the number of training tokens against its LLM Huggingface Benchmark score.